OpenStudy (anonymous):

Use Cramer's rule to solve the system. -4x + 3y + 3z = 12 -2y - 2z = -16 3x - 3z = -12

OpenStudy (anonymous):

\[\left[\begin{matrix}-4 & 3 & 3 & 12 \\ 0 & -2 & -2 & -16 \\ 3 & 0 & -3 & -12 \end{matrix}\right]\] so matrix looks like this, do you know the computation for cramers rule?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but with two, so how would it be with 3?

OpenStudy (anonymous):

\[x1 = \frac{\det(\left[\begin{matrix} 12 & 3 & 3 \\ -16 & -2 & -2 \\ -12 & 0 & -3 \end{matrix}\right])}{\det(\left[\begin{matrix} -4 & 3 & 3 \\ 0 & -2 & -2 \\ 3 & 0 & -3 \end{matrix}\right])}\] would be how to solve for x1 and you need to take the det of a 3x3, do you have a graphing calculator?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

then this is going to suck!!!!

OpenStudy (anonymous):

if i find a graphing calculator just plug in the numbers right?

OpenStudy (anonymous):

yes and then solve each matrix's determinant and divide them and get x1

OpenStudy (anonymous):

ok, thanks